# Large Language Models for Patient Comments Multi-Label Classification

### v1

@@ Abstract
Patient experience and care quality are crucial for a hospital’s sustainability and reputation. The analysis of patient feedback offers valuable insight into patient satisfaction and outcomes. However, the unstructured nature of these comments poses challenges for traditional machine learning methods following a supervised learning paradigm. This is due to the unavailability of labeled data and the nuances these texts encompass. This research explores leveraging Large Language Models (LLMs) in conducting Multi-label Text Classification (MLTC) of inpatient comments shared after a stay in the hospital. GPT-4o-Turbo was leveraged to conduct the classification. However, given the sensitive nature of patients’ comments, a security layer is introduced before feeding the data to the LLM through a Protected Health Information (PHI) detection framework, which ensures patients’ de-identification. Additionally, using the prompt engineering framework, zero-shot learning, in-context learning, and chain-of-thought prompting were experimented with. Results demonstrate that GPT-4o-Turbo, whether following a zero-shot or few-shot setting, outperforms traditional methods and Pre-trained Language Models (PLMs) and achieves the highest overall performance with an F1-score of 76.12% ± 0.021 and a weighted F1-score of 73.61% ± 0.006 followed closely by the few-shot learning results. Subsequently, the results’ association with other patient experience structured variables (e.g., rating) was conducted. The study enhances MLTC through the application of LLMs, offering healthcare practitioners an efficient method to gain deeper insights into patient feedback and deliver prompt, appropriate responses.
@@ 1 Introduction
  • PHI Detection Framework: Healthcare data is known for being sensitive and should be kept private. To de-identify the patients, a process that consists of text mining techniques (e.g., Regular Expressions) is developed to remove any PHI according to HIPAA. Additionally, all the LLMs involved are either used through their corresponding checkpoints or via API requests where data is not used for training, such as in the case of OpenAI’s GPT-4-Turbo (OpenAI, 2024a).
@@ 2 Related Work
Table 1: Table 1. Machine Learning Applications for Patient Comments Classification
@@ 3.3 LLM-based MLTC
MLTC, in the context of this paper, is a task consisting of assigning relevant topics to a given comment shared by a patient in the HCAHPS. Unlike single-label classification, which can be binary or multi-class, MLTC permits one feedback to be associated with one or more labels from a predefined set. This research explores the efficiency of using a high-performing LLM, GPT-4-Turbo, for this purpose. This LLM was developed by OpenAI (2024b) and is a cost-optimized and faster version of GPT-4. In addition to an advanced natural language understanding and generation, this LLM is able to handle various NLP tasks such as text classification and summarization. It is also a general-purpose LLM, which means that it was trained on various text domains. Prompt engineering was employed to stay within the optic of conducting a compute-efficient MLTC. A prompt was carefully crafted to include the predefined set of topics and their corresponding definitions, in addition to the classification instructions (similar to those shared with the group of annotators). Experiments using 0-shot learning and few-shot learning (1-shot, 3-shot, and 5-shot) were performed. In parallel, classifications using traditional machine learning methods relying on vectorization and multi-label transformation techniques as well as PLMs (i.e., BERT and BART) were also made. Their respective results were evaluated and contrasted using multiple classification metrics. For individual topic evaluation, F1-score and AUC-score were employed. Macro, micro, and weighted F1-scores were utilized to assess the overall MLTC performance. The MLTC was performed three times on the annotated patient comments samples.
@@ 4.1 Classification Evaluation
Table 4: Table 2. Traditional Machine Learning and Pre-trained Language Models Results
@@
Table 5: Table 3. GPT-4-Turbo In-Context Learning Results
@@
Table 3 summarizes the performance metrics for different few-shot learning scenarios using GPT-4-Turbo. The scenarios are evaluated on the ten predefined topics using F1 and AUC scores’ means and corresponding standard deviation resulting from the three different samples annotated for the case of 0-shot learning. They include 0-shot, where the LLM performs without providing specific comments classification examples, in addition to 1-shot, 3-shot, and 5-shot scenarios, where examples are included in the prompt for better guidance. The table shows that the performance varies depending on the topic and the scenario; however, despite providing examples in the prompt that led to the performance improvement of several cases, no consistency across all topics can be deduced. The 0-shot learning performs best in cases of topics like ‘Noisy Environment’, ‘Issues with Food Service’, ‘Room-related Issues’, and ‘Discharge-related Issues’, which are more straightforward and clear topics, neither subject to further interpretations nor encompassing multiple aspects. 1-shot learning also exhibits the highest F1-scores for multiple topics, including ‘Missing Belongings’, ‘Staff-related Issues’, and ‘Long Waiting Time’. Regarding the topics, ‘Positive Feedback,’ ‘Noisy Environment,’ ‘Missing Personal Belongings,’ ‘Issues with Food Service,’ and ’Room-related Issues’ demonstrate high F1-scores exceeding 80%. ‘Missing Personal Belongings’ shows an exceptional AUC score (>99%). In contrast, ‘Long Waiting Time’ and ‘Medical-related Issues’ resulted in relatively lower F1-scores than other topics. It is worth mentioning that experiments with Chain-of-Thought (COT) Prompting were also conducted, but since the results were not satisfactory, they were not reported.
Table 6: Table 5. Overall MLTC Results
@@
The results suggest that the LLM-based classifications of patient feedback are strong and significantly outperform both traditional machine learning and PLMs approaches. GPT-4-Turbo, especially in the 0-shot scenario, demonstrate that this general-purpose LLM has a strong inherent understanding of the patient comments context and is able to conduct an effective classification without neither further fine-tuning nor in-context learning especially that 0-shot learning performs interestingly well in the case of multiple topics, sometimes even outperforming few-shot learning. Furthermore, minimal variation across all four metrics can be remarked between the different few-shot scenarios which might indicate that the LLM’s base knowledge is sufficient for this MLTC and including examples in the prompt may just increase the number of input tokens without necessarily significantly ameliorating the performance.
@@ 4.2 Multiple-choice Questions’ Answers & Patient Demographics vs Classification Topics
An association analysis is conducted to understand how different classification topics expressed in patient experience feedback are related to the quantifiable HCAHPS responses and general information about the patients regarding their hospital stay and demographics. The analysis contrasts each one of the topics (i.e., a binary variable) that results from QUAD-LLM-MLTC classification with the collected quantifiable variables (i.e., categorical variables). The quantifiable variables are categorized into three sets: patient demographics and hospital stay-related, HCHAPS multiple-choice questions answers, and HCAHPS ratings.
@@
Table 7: Table 6. MLTC Outputs and Patients Demographics & Hospital Stay Association Analysis
@@
Table 8: Table 7. MLTC Outputs and HCAHPS Ratings Association Analysis
@@
Table 9: Table 8. MLTC Outputs and HCAHPS Multiple-Choice Questions Answers Association Analysis
@@ 4.3 Overall Rating vs Classification Topics
Table 10: Table 9. Regression & Correlation Analyses Summary

### v3

@@ Abstract
Patient experience and care quality are crucial for a hospital’s sustainability and reputation. The analysis of patient feedback offers valuable insight into patient satisfaction and outcomes. However, the unstructured nature of these comments poses challenges for traditional machine learning methods following a supervised learning paradigm. This is due to the unavailability of labeled data and the nuances these texts encompass. This research explores leveraging Large Language Models (LLMs) in conducting Multi-label Text Classification (MLTC) of inpatient comments shared after a stay in the hospital. GPT-4 Turbo was leveraged to conduct the classification. However, given the sensitive nature of patients’ comments, a security layer is introduced before feeding the data to the LLM through a Protected Health Information (PHI) detection framework, which ensures patients’ de-identification. Additionally, using the prompt engineering framework, zero-shot learning, in-context learning, and chain-of-thought prompting were experimented with. Results demonstrate that GPT-4 Turbo, whether following a zero-shot or few-shot setting, outperforms traditional methods and Pre-trained Language Models (PLMs) and achieves the highest overall performance with an F1-score of 76.12% ± 0.021 and a weighted F1-score of 73.61% ± 0.006 followed closely by the few-shot learning results. Subsequently, the results’ association with other patient experience structured variables (e.g., rating) was conducted. The study enhances MLTC through the application of LLMs, offering healthcare practitioners an efficient method to gain deeper insights into patient feedback and deliver prompt, appropriate responses.
@@ 1 Introduction
  • PHI Detection Framework: Healthcare data is known for being sensitive and should be kept private. To de-identify the patients, a process that consists of text mining techniques (e.g., Regular Expressions) is developed to remove any PHI according to HIPAA. Additionally, all the LLMs involved are either used through their corresponding checkpoints or via API requests where data is not used for training, such as in the case of OpenAI’s GPT-4 Turbo (OpenAI, 2024a).
@@ 2 Related Work
Table 1. Machine Learning Applications for Patient Comments Classification
@@ 3.3 LLM-based MLTC
MLTC, in the context of this paper, is a task consisting of assigning relevant topics to a given comment shared by a patient in the HCAHPS. Unlike single-label classification, which can be binary or multi-class, MLTC permits one feedback to be associated with one or more labels from a predefined set. This research explores the efficiency of using a high-performing LLM, GPT-4 Turbo, for this purpose. This LLM was developed by OpenAI (2024b) and is a cost-optimized and faster version of GPT-4. In addition to an advanced natural language understanding and generation, this LLM is able to handle various NLP tasks such as text classification and summarization. It is also a general-purpose LLM, which means that it was trained on various text domains. Prompt engineering was employed to stay within the optic of conducting a compute-efficient MLTC. A prompt was carefully crafted to include the predefined set of topics and their corresponding definitions, in addition to the classification instructions (similar to those shared with the group of annotators). Experiments using 0-shot learning and few-shot learning (1-shot, 3-shot, and 5-shot) were performed. In parallel, classifications using traditional machine learning methods relying on vectorization and multi-label transformation techniques as well as PLMs (i.e., BERT and BART) were also made. Their respective results were evaluated and contrasted using multiple classification metrics. For individual topic evaluation, F1-score and AUC-score were employed. Macro, micro, and weighted F1-scores were utilized to assess the overall MLTC performance. The MLTC was performed three times on the annotated patient comments samples.
@@ 4.1 Classification Evaluation
Table 2. Traditional Machine Learning and Pre-trained Language Models Results
@@
Table 3. GPT-4 Turbo In-Context Learning Results
@@
Table 3 summarizes the performance metrics for different few-shot learning scenarios using GPT-4 Turbo. The scenarios are evaluated on the ten predefined topics using F1 and AUC scores’ means and corresponding standard deviation resulting from the three different samples annotated for the case of 0-shot learning. They include 0-shot, where the LLM performs without providing specific comments classification examples, in addition to 1-shot, 3-shot, and 5-shot scenarios, where examples are included in the prompt for better guidance. The table shows that the performance varies depending on the topic and the scenario; however, despite providing examples in the prompt that led to the performance improvement of several cases, no consistency across all topics can be deduced. The 0-shot learning performs best in cases of topics like ‘Noisy Environment’, ‘Issues with Food Service’, ‘Room-related Issues’, and ‘Discharge-related Issues’, which are more straightforward and clear topics, neither subject to further interpretations nor encompassing multiple aspects. 1-shot learning also exhibits the highest F1-scores for multiple topics, including ‘Missing Belongings’, ‘Staff-related Issues’, and ‘Long Waiting Time’. Regarding the topics, ‘Positive Feedback,’ ‘Noisy Environment,’ ‘Missing Personal Belongings,’ ‘Issues with Food Service,’ and ’Room-related Issues’ demonstrate high F1-scores exceeding 80%. ‘Missing Personal Belongings’ shows an exceptional AUC score (>99%). In contrast, ‘Long Waiting Time’ and ‘Medical-related Issues’ resulted in relatively lower F1-scores than other topics. It is worth mentioning that experiments with Chain-of-Thought (COT) Prompting were also conducted, but since the results were not satisfactory, they were not reported.
Table 5. Overall MLTC Results
@@
The results suggest that the LLM-based classifications of patient feedback are strong and significantly outperform both traditional machine learning and PLMs approaches. GPT-4 Turbo, especially in the 0-shot scenario, demonstrate that this general-purpose LLM has a strong inherent understanding of the patient comments context and is able to conduct an effective classification without neither further fine-tuning nor in-context learning especially that 0-shot learning performs interestingly well in the case of multiple topics, sometimes even outperforming few-shot learning. Furthermore, minimal variation across all four metrics can be remarked between the different few-shot scenarios which might indicate that the LLM’s base knowledge is sufficient for this MLTC and including examples in the prompt may just increase the number of input tokens without necessarily significantly ameliorating the performance.
@@ 4.2 Multiple-choice Questions’ Answers & Patient Demographics vs Classification Topics
An association analysis is conducted to understand how different classification topics expressed in patient experience feedback are related to the quantifiable HCAHPS responses and general information about the patients regarding their hospital stay and demographics. The analysis contrasts each one of the topics (i.e., a binary variable) that results from the text classification with the collected quantifiable variables (i.e., categorical variables). The quantifiable variables are categorized into three sets: patient demographics and hospital stay-related, HCHAPS multiple-choice questions answers, and HCAHPS ratings.
@@
Table 6. MLTC Outputs and Patients Demographics & Hospital Stay Association Analysis
@@
Table 7. MLTC Outputs and HCAHPS Ratings Association Analysis
@@
Table 8. MLTC Outputs and HCAHPS Multiple-Choice Questions Answers Association Analysis
@@ 4.3 Overall Rating vs Classification Topics
Table 9. Regression & Correlation Analyses Summary
@@ Acknowledgement
[1] Doyle, C., Lennox, L., & Bell, D. (2013). A systematic review of evidence on the links between patient experience and clinical safety and effectiveness. BMJ Open, 3(1).
[2] Boulding, W., Glickman, S. W., Manary, M. P., Schulman, K. A., & Staelin, R. (2011). Relationship between patient satisfaction with inpatient care and hospital readmission within 30 days. The American Journal of Managed Care, 17(1), 41-48.
[3] Senot, C., Chandrasekaran, A., & Ward, P. T. (2016). The impact of combining conformance and experiential quality on hospitals’ readmissions and cost performance. Management Science, 62(3), 829-848.
[4] Meterko, M., Wright, S., Lin, H., Lowy, E., & Cleary, P. D. (2010). Mortality among patients with acute myocardial infarction: the influences of patient-centered care and evidence-based medicine. Health Services Research, 45(5p1), 1188-1204.
[5] Shenoy, A. (2021). Patient safety from the perspective of quality management frameworks: a review. Patient Safety in Surgery, 15(1), 12.
[6] Centers for Medicare & Medicaid Services (CMS). (2018). Hospital Value-Based Purchasing. Retrieved from CMS website.
[7] Greaves, F., Ramirez-Cano, D., Millett, C., Darzi, A., & Donaldson, L. (2013). Use of sentiment analysis for capturing patient experience from free-text comments posted online. Journal of Medical Internet Research, 15(11), e239.
[8] Manary, M. P., Boulding, W., Staelin, R., & Glickman, S. W. (2013). The patient experience and health outcomes. The New England Journal of Medicine, 368(3), 201-203.
[9] Gallan, A. S., Niraj, R., & Singh, A. (2022). Beyond HCAHPS: Analysis of patients’ comments provides an expanded view of their hospital experiences. Patient Experience Journal, 9(1), 159-168.
[10] Shovkun, M. (2018). Manual and automatic analysis of patients values and preferences using Seton HCAHPS surveys (Doctoral dissertation).
[11] Huppertz, J. W., & Otto, P. (2018). Predicting HCAHPS scores from hospitals’ social media pages: A sentiment analysis. Health care management review, 43(4), 359-367.
[12] Huppertz, J. W., & Smith, R. (2014). The value of patients’ handwritten comments on HCAHPS surveys. Journal of Healthcare Management, 59(1), 31-47.
[13] Paul, R., Pandit, A., & Bhardwaj, R. (2022). Transforming Healthcare through Sentiment Analysis: Tool for Patient Satisfaction. Journal of Algebraic Statistics, 13(3), 3962-3980.
[14] OpenAI. (OpenAI, 2024a). Enterprise Privacy. Retrieved June 8, 2024, from https://openai.com/enterprise-privacy/ [15] Mera, K., & Ichimura, T. (2008). User’s Comment Classifying Method Using Self Organizing Feature Map on Healthcare System for Diabetic. In 4th International Workshop on Computational Intelligence & Applications Proceedings: IWCIA 2008 (pp. 31-36). IEEE SMC Hiroshima Chapter.
[16] Alemi, F., Torii, M., Clementz, L., & Aron, D. C. (2012). Feasibility of real-time satisfaction surveys through automated analysis of patients’ unstructured comments and sentiments. Quality Management in Healthcare, 21(1), 9-19.
[17] Wagland, R., Recio-Saucedo, A., Simon, M., Bracher, M., Hunt, K., Foster, C., … & Corner, J. (2016). Development and testing of a text-mining approach to analyse patients’ comments on their experiences of colorectal cancer care. BMJ quality & safety, 25(8), 604-614.
[18] Pan, Q., Li, H., Chen, D., & Sun, K. (2018, June). Sentiment analysis of medical comments based on character vector convolutional neural networks. In 2018 IEEE Symposium on Computers and Communications (ISCC) (pp. 1-4). IEEE.
[19] Shah, A. M., Yan, X., Shah, S. J., & Khan, S. (2018). Use of sentiment mining and online NMF for topic modeling through the analysis of patients online unstructured comments. In Smart Health: International Conference, ICSH 2018, Wuhan, China, July 1–3, 2018, Proceedings 6 (pp. 191-203). Springer International Publishing.
[20] Hussain, S., Nasir, A., Aslam, K., Tariq, S., & Ullah, M. F. (2020). Predicting mental illness using social media posts and comments. International Journal of Advanced Computer Science and Applications, 11(12).
[21] Shan, Y., Zhong, Z., Che, C., Jin, B., & Wei, X. (2021, December). Aspect-level sentiment classification of chinese patient comments based on pre-trained sentiment embedding. In 2021 IEEE International Conference on Bioinformatics and Biomedicine (BIBM) (pp. 1503-1508). IEEE.
[22] Asghari, M., Nielsen, J., Gentili, M., Koizumi, N., & Elmaghraby, A. (2022). Classifying comments on social media related to living kidney donation: Machine learning training and validation study. JMIR Medical Informatics, 10(11), e37884.
[23] Sakai, H., Mikaeili, M., Lam, S. S., & Bosire, J. (2023). Text Classification for Patient Experience Improvement: A Neural Network Approach. In IIE Annual Conference. Proceedings (pp. 1-6). Institute of Industrial and Systems Engineers (IISE).
[24] Linton, A. G., Dimitrova, V., Downing, A., Wagland, R., & Glaser, A. (2023). Weakly Supervised Text Classification on Free Text Comments in Patient-Reported Outcome Measures. arXiv preprint arXiv:2308.06199.
[25] Alhazzani, N. Z., Al-Turaiki, I. M., & Alkhodair, S. A. (2023). Text Classification of Patient Experience Comments in Saudi Dialect Using Deep Learning Techniques. Applied Sciences, 13(18), 10305.
[26] Sakai, H., Lam, S. S., Mikaeili, M., & Bosire, J. (2024). Patient Experience Feedback Sentiment Analysis: Combining BERT and LSTM with Genetic Algorithm Optimization. In IISE Annual Conference. Proceedings (pp. 1-6). Institute of Industrial and Systems Engineers (IISE).
[27] Mæhlum, P., Samuel, D., Norman, R. M., Jelin, E., Bjertnæs, Ø. A., Øvrelid, L., & Velldal, E. (2024). It’s Difficult to be Neutral–Human and LLM-based Sentiment Annotation of Patient Comments. arXiv preprint arXiv:2404.18832.
[28] Babbar, R., & Schölkopf, B. (2017, February). Dismec: Distributed sparse machines for extreme multi-label classification. In Proceedings of the tenth ACM international conference on web search and data mining (pp. 721-729).
[29] Bhatia, K., Jain, H., Kar, P., Varma, M., & Jain, P. (2015). Sparse local embeddings for extreme multi-label classification. Advances in neural information processing systems, 28.
[30] Kong, H. J. (2019). Managing unstructured big data in healthcare system. Healthcare informatics research, 25(1), 1.
[31] HIPAA Journal. (2024). What is considered PHI under HIPAA? HIPAA Journal. Retrieved from https://www.hipaajournal.com/considered-phi-hipaa/
[32] OpenAI. (OpenAI, 2024b). GPT-4 Turbo in the OpenAI API. OpenAI Help Center. Retrieved from https://help.openai.com/en/articles/8555510-gpt-4-turbo-in-the-openai-ap
[33] Cohen, J. (2013). Statistical power analysis for the behavioral sciences. routledge.
[34] Kim, H. Y. (2017). Statistical notes for clinical researchers: Chi-squared test and Fisher’s exact test. Restorative dentistry & endodontics, 42(2), 152-155.
[35] Mikaeili, M., Lam, S. S., & Bosire, J. (2020). Text-Mining Framework for Hospital Patient Experience Surveys: A Case Study. In IIE Annual Conference. Proceedings (pp. 1-6). Institute of Industrial and Systems Engineers (IISE).
6 Appendix
6.1 Topics Details
Table 11. Topics Identified and Their Detailed Descriptions
Topic Details
[…]